What is the solution set of:
x - 4y = 8
y = -5x + 7
Thank you!

Answer by thesvw(77)

x - 4y = 8
y = -5x + 7
Substitute y to the first equation.
x -4(-5x+7) =8
x +20x -28 =8
21x = 36
x= 36/21
Substitute this x value to second equation to find the y value.